# Log-majorizations between quasi-geometric type means for matrices

### v2

@@ 3 Log-majorizations
Let us begin with surveying the log-majorizations known so far between quasi-geometric type means $R_{\alpha,p}$, $G_{\alpha,p}$ and $SG_{\alpha,p}$ in the following theorem.
@@
Indeed, (a) and (b) are the log-majorizations due to Araki [4] and Ando and Hiai [3, Theorem 2.1]), respectively. In [39] the matrix perspective $G_{\beta}(A,B):=A^{1/2}(A^{-1/2}BA^{-1/2})^{\beta}A^{1/2}$ for $1\leq\beta<0$ was treated, where $G_{\beta}(A,B)$ is denoted by $A\natural_{\beta}B$. Since $G_{\beta}(A,B)=G_{1-\beta}(B,A)$ and $1<1-\beta\leq 2$, the log-majorization in (c) is equivalent to that shown by Kian and Seo [39, Theorem 3.1]. (d) was shown in [31, Corollary 5.2], and (e) and (f) were shown in [31, Proposition 5.1(a), (b)]. Finally, (g) is a rewriting of the log-majorization due to Gan and Tam [24, Theorem 3.7]. By Proposition 2.1(2) note that the log-majorizations in the $0<\alpha<1$ case of (a), (b) and (e) hold for all $A,B\geq 0$ and those in the $\alpha>1$ case of (a), (c), (d), (f) and (g) hold for $A,B\geq 0$ with $s(A)\geq s(B)$.
@@ 3.1 $SG_{\alpha,p}\prec_{\log}R_{\alpha,q}$ and $R_{\alpha,q}\prec_{\log}SG_{\alpha,p}$
First we recall that the log-majorizations in (e) and (f) of Theorem 3.1 were indeed presented in [31] in the ‘if and only if’ statement. For the convenience to make explicit comparison with the log-majorizations shown in Section 3 below, we include [31, Proposition 5.1] in its complete form in the following:
@@
When $0<\alpha<1$, the next theorem gives the necessary and sufficient condition on $p,q$ for $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ (resp., $R_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$) to hold for all $A,B\geq 0$ with $s(A)\geq s(B)$, thus strengthening the assertion in (g) above.
@@
where $r:=q/p$. Conversely, for any $X,Y>0$, setting $A:=Y^{1/p}$ and $B:=(XYX)^{1/p}$ we have $Y=A^{p}$ and $X=A^{-p}\#B^{p}$. Furthermore, note that for $X_{1},X_{2}\in\mathbb{M}_{2}^{++}$ with $\det X_{1}=\det X_{2}$, $X_{1}\prec_{\log}X_{2}$ if and only if $\lambda_{1}(X_{1})\leq\lambda(X_{2})$. Hence we see that condition (ii) is equivalent to saying that
@@
Now let $X:=A_{0}$ and $Y:=B_{\theta}$ in (2.10) for $x,y>0$ with $x^{2}y\neq 1$ and $x^{2\alpha}y\neq 1$, and argue as in the proof of [33, Theorem 4.23]. Since
@@
one can apply [33, Example 3.2(1)] to compute
@@
When $x^{2\alpha}y<1$, we apply [33, Lemma 3.6] to (3.4) and (3.3) to obtain
@@
Letting $y\searrow 0$ gives $r\leq-1-\alpha r$ so that $1\leq(1-\alpha)r$, i.e., $p/q\leq 1-\alpha$. Since $R_{\alpha,p}(A,B)=R_{1-\alpha,p}(B,A)$ and $SG_{\alpha,p}(A,B)=SG_{1-\alpha,p}(A,B)$ (see [41, Proposition 4.2(iii)]) for all $A,B>0$, note that (ii) implies condition (ii) with $1-\alpha$ in place of $\alpha$. Hence we have $p/q\leq\alpha$ too so that (iii) follows.
@@
When $\alpha>1$, no necessary condition on $p,q>0$ under which $SG_{\alpha,p}\prec_{\log}R_{\alpha,q}$ holds is known. In fact, when $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ holds for all $A,B\in\mathbb{M}_{2}^{++}$, one has the reversed inequality of (3.9) but nothing follows from that. This suggests us that $SG_{\alpha,p}\prec_{\log}R_{\alpha,q}$ might hold for any $p,q>0$.
@@
Now let $Y:=A_{0}$ and $X:=B_{\theta}$ in (2.10) for $x,y>0$ with $xy\neq 1$, $x^{1-\alpha}y\neq 1$ and $x^{2(1-\alpha)}y\neq 1$. From the proof of [33, Theorem 4.35] we have
@@
Hence by [33, Example 3.2(1)] we compute
@@
When $x^{1-\alpha}y<1$ and $x^{2(1-\alpha)y}<1$, applying [33, Lemma 3.6] to (3.11) and (3.14) we have
@@
which is exactly the BLP log-majorization due to Bebiano–Lemos–Providência [6, Theorem 2.1].
@@
Let $\alpha\in(0,\infty)\setminus\{1\}$ and $p,q>0$. For any pair $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ as mentioned in Remark 2.4, note that $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}\mathcal{N}_{\alpha,q}(A,B)$ holds for all $A,B\in\mathbb{M}_{2}^{++}$ if and only if $\lambda_{1}(\mathcal{M}_{\alpha,p}(A,B))\leq\lambda_{1}(\mathcal{N}_{\alpha,q}(A,B))$ for all $A,B\in\mathbb{M}_{2}^{++}$. Therefore, one can easily find a necessary condition for $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ by use of expressions given in Lemma 2.5. For instance, if $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$, then from (2.11) and (2.13) one must have
@@
While the log-majorizations $G_{\alpha,p}\prec_{\log}G_{\alpha,q}$ for $\alpha\in(0,2]\setminus\{1\}$ were completely characterized as stated in Theorem 3.1(b) and (c), the characterizations of $SG_{\alpha,p}\prec_{\log}SG_{\alpha,q}$ and $\widetilde{S}G_{\alpha,p}\prec_{\log}\widetilde{S}G_{\alpha,q}$, as well as $G_{\alpha,p}\prec_{\log}G_{\alpha,q}$ for $\alpha>2$, have not been obtained yet. In this subsection we discuss the problem though without complete success.
@@
    Assume that $\widetilde{S}G_{\alpha,p}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,q}(A,B)$ holds for all $A,B\in\mathbb{M}_{2}^{++}$. Then we have $p\leq q$ if $0<\alpha<1$, and $p\geq q$ if $\alpha>1$.
@@
What we are most interested in is whether a necessary condition $p\leq q$ or $p\geq q$ confirmed in Proposition 3.11 is indeed sufficient for $SG_{\alpha,p}\prec_{\log}SG_{\alpha,q}$ and $\widetilde{S}G_{\alpha,p}\prec_{\log}\widetilde{S}G_{\alpha,q}$ to hold or not. The problem is to show the variants of Ando–Hiai’s inequality [3] for $SG_{\alpha,p}$ and $\widetilde{S}G_{\alpha,q}$. Note that the condition on $p,q$ in Proposition 3.11(1) and that in Proposition 3.11(2) and (3) are opposite. This suggests us that two types of weighted spectral geometric means are quite different from Kubo–Ando’s weighted geometric mean $_{\alpha}$.
@@ 3.4 More log-majorizations
where the additional inequalities in both sides are seen from [33, Propositions 4.9, 4.19 and Remark 2.7(3)].
@@
    For any $0<\alpha<1$ and $p>0$ there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $\widetilde{S}G_{\alpha,p}(A,B)\not\prec_{\log}LE_{\alpha}(A,B)$.
@@
(1) follows by letting $p\searrow 0$ in (c) and (d) of Theorem 3.1. (2) and (3) are immediately seen by comparing (2.15) with (2.13) and (2.14), respectively. (4) and (5) follow as the $q\searrow 0$ limits (due to Theorem 2.2) of the log-majorizations in Theorems 3.6(2) and 3.7(2), respectively. ∎
@@
On the other hand, apply [33, Example 3.2(3)] to (3.12) and (3.13) to compute
@@ 4 Equality cases in norm inequalities derived from log-majorizations
Recall that if $X,Y\in\mathbb{M}_{n}^{+}$ and $X\prec_{\log}Y$, then $\|X\|\leq\|Y\|$ holds for all unitarily invariant norms $\|\cdot\|$ on $\mathbb{M}_{n}$; see, e.g., [29, Proposition 4.4.13]. Therefore, if a pair $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,p})$ of quasi-geometric type means satisfies $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$, then we have $\|\mathcal{M}_{\alpha,p}(A,B)\|\leq\|\mathcal{N}_{\alpha,q}(A,B)\|$ for all $A,B\geq 0$ and for any unitarily invariant norm $\|\cdot\|$. In this section we aim at characterizing the equality case in the above norm inequality in terms of commutativity $AB=BA$ in certain cases of the pair $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,p})$ and the norm $\|\cdot\|$. This is reasonable since $AB=BA$ implies that $\mathcal{M}_{\alpha,p}(A,B)=\mathcal{N}_{\alpha,q}(A,B)$ holds for any pair of quasi-geometric type means. We say that a norm $\|\cdot\|$ on $\mathbb{M}_{n}$ is strictly increasing if $A\geq B\geq 0$ and $\|A\|=\|B\|$ imply $A=B$. The Schatten $p$-norms where $1\leq p<\infty$ (particularly, the trace norm) are typical examples of strictly increasing unitarily invariant norms.
@@
    Let $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ be any of (4.1)–(4.5) and $A,B\geq 0$, with an additional assumption $s(A)\geq s(B)$ for the $\alpha>1$ case of (4.1) and (4.2) and for (4.3) and (4.3). If $\|\mathcal{M}_{\alpha,p}(A,B)\|=\|\mathcal{N}_{\alpha,q}(A,B)\|$, then $AB=BA$
@@
(1) The main of this part is the assertion for (4.1), which was formerly shown in [28, Theorem 2.1] and the proof has been updated in [32, Appendix A]. [^2] We note that the claim [32, (A.3)] and its reasoning in the last five lines of [32, p. 544] is incorrect. But (A.3) is unnecessary in the proof of [32, Appendix]. Therefore, these lines and Remark A.2 of [32] should be deleted. The other cases in (4.2)–(4.5) are all reduced to (4.1). Indeed, assume that $\|G_{\alpha,p}(A,B)\|=\|R_{\alpha,q}(A,B)\|$ for $\alpha,p,q$ satisfying one of the three conditions in (4.2). When either $0<\alpha<1$ with $p,q>0$ arbitrary, or $\alpha>1$ and $p/q<\min\{\alpha/2,\alpha-1\}$, choose a $q^{\prime}\in(0,q)$ such that $\alpha,p,q^{\prime}$ satisfy the same condition. By Theorem 3.1(a) and (e) one has
@@
(2) The main of this part is the assertion for (4.6), which was formerly shown for $0<\alpha<1$ in [28, Theorem 3.1]. The assertion for $1<\alpha\leq 2$ is new. We defer its details to Appendix A because it seems more instructive to present them independently. The other cases in (4.7)–(4.9) are all reduced to (4.6). For instance, assume that $\|SG_{\alpha,p}(A,B)\|=\|G_{\alpha,q}(A,B)\|$ for $\alpha,p,q$ satisfying one of the two conditions in (4.7). When $0<\alpha<1$ with $p,q>0$ arbitrary, for a $q^{\prime}\in(0,q)$ by Theorem 3.1(b) and Proposition 3.13(1) one has
@@
It is not possible to have the local characterization similar to Proposition 4.2 for log-majorizaion between $G_{\alpha,p}$ and $R_{\alpha,q}$ for $\alpha>1$. In fact, log-majorization between $G_{\alpha,p}$ and $R_{\alpha,q}$ is indefinite when $\alpha>1$ and $\min\{\alpha/2,\alpha-1\}<p/q<\max\{\alpha/2,\alpha-1\}$; see Theorem 3.1(e) and (f).

### v3

@@ 3 Log-majorizations
Let us begin by surveying the log-majorizations known so far between quasi-geometric type means $R_{\alpha,p}$, $G_{\alpha,p}$ and $SG_{\alpha,p}$ in the following theorem.
@@
  • (h)
    $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ either if $1<\alpha\leq 3/2$ and $p=q$, or if $\alpha\in(0,1/2]\cup(1,2]$ and $p/q=\alpha$.
Indeed, (a) and (b) are the log-majorizations due to Araki [4] and Ando and Hiai [3, Theorem 2.1]), respectively. In [41] the matrix perspective $G_{\beta}(A,B):=A^{1/2}(A^{-1/2}BA^{-1/2})^{\beta}A^{1/2}$ for $1\leq\beta<0$ was treated, where $G_{\beta}(A,B)$ is denoted by $A\natural_{\beta}B$. Since $G_{\beta}(A,B)=G_{1-\beta}(B,A)$ and $1<1-\beta\leq 2$, the log-majorization in (c) is equivalent to that shown by Kian and Seo [41, Theorem 3.1]. (d) was shown in [33, Corollary 5.2], and (e) and (f) were shown in [33, Proposition 5.1(a), (b)]. (g) is a rewriting of the log-majorization due to Gan and Tam [26, Theorem 3.7]. Finally, (h) is rewriting of the log-majorizations shown by Furuichi and Seo [21, Theorems 4.3 and 4.8]. By Proposition 2.1(2) note that the log-majorizations in the $0<\alpha<1$ case of (a), (b), (e) hold for all $A,B\geq 0$, and those in the $\alpha>1$ case of (a), (c), (d), (e), (f) as well as those in (g), (h) hold for $A,B\geq 0$ with $s(A)\geq s(B)$.
@@ 3.1 $SG_{\alpha,p}\prec_{\log}R_{\alpha,q}$ and $R_{\alpha,q}\prec_{\log}SG_{\alpha,p}$
First we recall that the log-majorizations in (e) and (f) of Theorem 3.1 were indeed presented in [33] in the ‘if and only if’ statement. For the convenience to make explicit comparison with the log-majorizations shown in Section 3 below, we include [33, Proposition 5.1] in its complete form in the following:
@@
When $0<\alpha<1$, the next theorem gives the necessary and sufficient condition on $p,q$ for $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ (resp., $R_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$) to hold for all $A,B\geq 0$ with $s(A)\geq s(B)$, thus strengthening the assertion in (h) and (g) above.
@@
where $r:=q/p$. Conversely, for any $X,Y>0$, setting $A:=Y^{1/p}$ and $B:=(XYX)^{1/p}$ we have $Y=A^{p}$ and $X=A^{-p}\#B^{p}$. Hence in view of Remark 2.4(4) we see that condition (ii) is equivalent to saying that
@@
Now let $X:=A_{0}$ and $Y:=B_{\theta}$ in (2.10) for $x,y>0$ with $x^{2}y\neq 1$ and $x^{2\alpha}y\neq 1$, and argue as in the proof of [35, Theorem 4.23]. Since
@@
one can apply [35, Example 3.2(1)] to compute
@@
When $x^{2\alpha}y<1$, we apply [35, Lemma 3.6] to (3.4) and (3.3) to obtain
@@
Letting $y\searrow 0$ gives $r\leq-1-\alpha r$ so that $1\leq(1-\alpha)r$, i.e., $p/q\leq 1-\alpha$. Since $R_{\alpha,p}(A,B)=R_{1-\alpha,p}(B,A)$ and $SG_{\alpha,p}(A,B)=SG_{1-\alpha,p}(A,B)$ (see [44, Proposition 4.2(iii)]) for all $A,B>0$, note that (ii) implies condition (ii) with $1-\alpha$ in place of $\alpha$. Hence we have $p/q\leq\alpha$ too so that (iii) follows.
@@
When $\alpha>1$, no necessary condition on $p,q>0$ for $SG_{\alpha,p}\prec_{\log}R_{\alpha,q}$ to hold is known. In fact, when $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ holds for all $A,B\in\mathbb{M}_{2}^{++}$, one has the reversed inequality of (3.9) but nothing follows from that. This suggests us that $SG_{\alpha,p}\prec_{\log}R_{\alpha,q}$ might hold for any $p,q>0$.
@@
Now let $Y:=A_{0}$ and $X:=B_{\theta}$ in (2.10) for $x,y>0$ with $xy\neq 1$, $x^{1-\alpha}y\neq 1$ and $x^{2(1-\alpha)}y\neq 1$. From the proof of [35, Theorem 4.35] we have
@@
Hence by [35, Example 3.2(1)] we compute
@@
When $x^{1-\alpha}y<1$ and $x^{2(1-\alpha)y}<1$, applying [35, Lemma 3.6] to (3.11) and (3.14) we have
@@
which is exactly the BLP log-majorization due to Bebiano, Lemos and Providência [6, Theorem 2.1].
@@
Let $\alpha\in(0,\infty)\setminus\{1\}$ and $p,q>0$. For any pair $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ such as in Remark 2.4, one can easily find a necessary condition for $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ by use of expressions given in Lemma 2.5. For instance, if $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$, then from (2.11) and (2.13) one must have
@@
While the log-majorizations $G_{\alpha,p}\prec_{\log}G_{\alpha,q}$ for $\alpha\in(0,2]\setminus\{1\}$ were completely characterized as stated in Theorem 3.1(b) and (c), the characterizations of $SG_{\alpha,p}\prec_{\log}SG_{\alpha,q}$ and $\widetilde{S}G_{\alpha,p}\prec_{\log}\widetilde{S}G_{\alpha,q}$, as well as $G_{\alpha,p}\prec_{\log}G_{\alpha,q}$ for $\alpha>2$, have not been obtained yet. [^2] It was claimed in [26, Theorem 3.3] that $SG_{\alpha,p}(A,B)\prec_{\log}SG_{\alpha,q}(A,B)$ for all $A,B>0$ if $0<\alpha<1$ and $0<p\leq q$; however the proof contains a flaw. In this subsection we discuss the problem though without complete success.
@@
    Assume that $\widetilde{S}G_{\alpha,p}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,q}(A,B)$ holds for all $A,B\in\mathbb{M}_{2}^{++}$. Then we have $p\leq q$ whichever $0<\alpha<1$ or $\alpha>1$.
@@
What we are most interested in is whether a necessary condition $p\leq q$ or $p\geq q$ confirmed in Proposition 3.11 is indeed sufficient for $SG_{\alpha,p}\prec_{\log}SG_{\alpha,q}$ and $\widetilde{S}G_{\alpha,p}\prec_{\log}\widetilde{S}G_{\alpha,q}$ for relevant $\alpha$ to hold or not. The problem is to show the variants of Ando–Hiai’s inequality [3] for $SG_{\alpha,p}$ and $\widetilde{S}G_{\alpha,q}$. Regarding this, see remark (3) of Section 6 as well.
Note that for $0<\alpha<1$ the condition on $p,q$ in Proposition 3.11(1) and that in Proposition 3.11(2) and (3) are opposite. This suggests us that two types of weighted spectral geometric means are quite different from Kubo–Ando’s weighted geometric mean $_{\alpha}$.
@@ 3.4 More log-majorizations
where the additional inequalities in both sides are seen from [35, Propositions 4.9, 4.19 and Remark 2.7(3)].
@@
    For any $\alpha\in(0,1)$ and $p>0$ there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $SG_{\alpha,p}(A,B)\not\prec_{\log}LE_{\alpha}(A,B)$.
  • (3)
@@
    For any $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$ there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $\widetilde{S}G_{\alpha,p}(A,B)\not\prec_{\log}LE_{\alpha}(A,B)$.
@@
(1) follows by letting $p\searrow 0$ in (c) and (d) of Theorem 3.1. (2) and (3) are immediately seen by comparing (2.15) with (2.13), and (4) is by comparing (2.15) with (2.14). (5) and (6) follow as the $q\searrow 0$ limits (due to Theorem 2.2) of the log-majorizations in Theorems 3.6(2) and 3.7(2), respectively. ∎
@@
On the other hand, apply [35, Example 3.2(3)] to (3.12) and (3.13) to compute
@@ 4 Equality cases in norm inequalities derived from log-majorizations
Recall that if $X,Y\in\mathbb{M}_{n}^{+}$ and $X\prec_{\log}Y$, then $\|X\|\leq\|Y\|$ holds for all unitarily invariant norms $\|\cdot\|$ on $\mathbb{M}_{n}$; see, e.g., [31, Proposition 4.4.13]. Therefore, if a pair $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,p})$ of quasi-geometric type means satisfies $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$, then we have $\|\mathcal{M}_{\alpha,p}(A,B)\|\leq\|\mathcal{N}_{\alpha,q}(A,B)\|$ for all $A,B\geq 0$ (with $s(A)\geq s(B)$) and for any unitarily invariant norm $\|\cdot\|$. In this section we aim at characterizing the equality case in the above norm inequality in terms of commutativity $AB=BA$ in certain cases of the pair $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,p})$ and the norm $\|\cdot\|$. This is reasonable since $AB=BA$ implies that $\mathcal{M}_{\alpha,p}(A,B)=\mathcal{N}_{\alpha,q}(A,B)$ holds for any pair of quasi-geometric type means. We say that a norm $\|\cdot\|$ on $\mathbb{M}_{n}$ is strictly increasing if $A\geq B\geq 0$ and $\|A\|=\|B\|$ imply $A=B$. The Schatten $p$-norms where $1\leq p<\infty$ (particularly, the trace norm) are typical examples of strictly increasing unitarily invariant norms.
@@
    Let $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ be any of (4.1)–(4.5) and $A,B\geq 0$, with an additional assumption $s(A)\geq s(B)$ for the $\alpha>1$ case of (4.1), (4.2) and (4.5), and for (4.3) and (4.4). If $\|\mathcal{M}_{\alpha,p}(A,B)\|=\|\mathcal{N}_{\alpha,q}(A,B)\|$, then $AB=BA$
@@
(1) The main of this part is the assertion for (4.1), which was formerly shown in [30, Theorem 2.1] and the proof has been updated in [34, Appendix A]. [^3] We note that the claim [34, (A.3)] and its reasoning in the last five lines of [34, p. 544] are incorrect. But (A.3) is unnecessary in the proof of [34, Appendix]. Therefore, these lines and Remark A.2 of [34] should be deleted. The other cases in (4.2)–(4.5) are all reduced to (4.1). Indeed, assume that $\|G_{\alpha,p}(A,B)\|=\|R_{\alpha,q}(A,B)\|$ for $\alpha,p,q$ satisfying one of the three conditions in (4.2). When either $0<\alpha<1$ with $p,q>0$ arbitrary, or $\alpha>1$ and $p/q<\min\{\alpha/2,\alpha-1\}$, choose a $q^{\prime}\in(0,q)$ such that $\alpha,p,q^{\prime}$ satisfy the same condition. By Theorem 3.1(a) and (e) one has
@@
(2) The main of this part is the assertion for (4.6), which was formerly shown for $0<\alpha<1$ in [30, Theorem 3.1]. The assertion for $1<\alpha\leq 2$ is new. We defer its details to Appendix A because it seems more instructive to present them independently. The other cases in (4.7)–(4.9) are all reduced to (4.6). For instance, assume that $\|SG_{\alpha,p}(A,B)\|=\|G_{\alpha,q}(A,B)\|$ for $\alpha,p,q$ satisfying one of the two conditions in (4.7). When $0<\alpha<1$ with $p,q>0$ arbitrary, for a $q^{\prime}\in(0,q)$ by Theorem 3.1(b) and Proposition 3.13(1) one has
@@
It is not possible to have the local characterization similar to Proposition 4.2 for log-majorizaions between $G_{\alpha,p}$ and $R_{\alpha,q}$ for $\alpha>1$. In fact, log-majorizations between $G_{\alpha,p}$ and $R_{\alpha,q}$ are indefinite when $\alpha>1$ and $\min\{\alpha/2,\alpha-1\}<p/q<\max\{\alpha/2,\alpha-1\}$; see Theorem 3.1(e) and (f).